# CrossDehaze: Scaling Up Image Dehazing with Cross-Data Vision Alignment and Augmentation

### v1

@@ V Conclusion
In this paper, we propose a novel method that performs data augmentation separately on the external and internal aspects of the data. Specifically, we achieve external data augmentation using channel transfer and internal data augmentation using strong-weak augmentation self-supervised learning. Our method leverages data diversity at the external level and further explores internal information to effectively dehaze images. Experimental results show that our method achieves outstanding performance after applying both internal and external data augmentation methods. Compared to other advanced methods, our method performs better in terms of PSNR and SSIM metrics on both the Natural Image Dataset (NID) and the Remote Sensing Image Dataset (RSID). The resulting dehazed images also contain rich details and closely resemble the ground truth haze-free images.

### v2

@@ V Conclusion and Limitation
In this paper, we propose a novel method that performs data augmentation separately on the external and internal aspects of the data. Specifically, we achieve external- augmentation and internal- augmentation using weak-to-strong self-supervised learning. Our method leverages data diversity at the external level and further explores internal information to effectively dehaze images.
Although our method achieves superior results in both quantitative and qualitative evaluations, the introduction of augmentation operations has led to an increase in training time compared to other methods. In the future, we will focus on research aimed at improving training efficiency.